The pure rotational spectrum of 1H131I consists of a series of lines separated by 13.10 cm−1

Calculate


(a) the moment of interia, and


(b) the internuclear distance.



1
Expert's answer
2021-03-26T05:43:13-0400

Moment of inertia = r^2

= 13.1^2

= 171.61 kg. cm^2

Internuclear distance = 171.61/131×13.1 × 10^1

= 171.61pm